OpenStudy (anonymous):

The governor of state A earns $54,215 more than the governor of state B. if the total of their salaries is $289,675 find the salaries of each.

OpenStudy (anonymous):

total is \(A+B=289,675\) and you know \(A=B+54,215\) so \[2B+54,215=289,675\] solve for \(B\)

OpenStudy (anonymous):

you good from there? subtract \(54,215\) from both sides, get \[2B=235,460\] then divide by 2 to get \(B\)

OpenStudy (anonymous):

oh ok right i get it

OpenStudy (anonymous):

good!

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

yw